OpenStudy (anonymous):

The equation e^(4x) + x=2 has a solution near x=0. By replacing the left side of the equation by its linearization, find an approximate value for the solution. **Write the EXACT answer** x=_______ please explain!!! Thank you:)

OpenStudy (anonymous):

 in case it wasn't clear when typed:)

OpenStudy (freckles):

the linearization method find f'(x) if f(x)=e^(4x)+x then the linear equation that f(x) is approximately y=mx+b where m=f'(0) in case since we want to know what happens near x=0 the y-intercept can be found by using a point on the line which we know to be (0,f(0))=(0,1) correct?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

so we need to find f'(x) first?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

so would f'(x)= e^(4x) + 0 = e^(4x)?

OpenStudy (freckles):

f'(x)=(4x)'e^(4x)+(x)'

OpenStudy (freckles):

(4x)'=4 (x)'=1

OpenStudy (anonymous):

ohh darn... using the chain rule?

OpenStudy (freckles):

Yippie

OpenStudy (anonymous):

so f'(x)= 5 ?

OpenStudy (freckles):

So we have f'(0) is totally 5 good job so we have y=5x+b we need the y-intercept now

OpenStudy (freckles):

we know this line goes through (0,e^(4*0)+0)

OpenStudy (freckles):

Use that to find b

OpenStudy (anonymous):

so 0=5(0)+b b=0?

OpenStudy (freckles):

e^0=?

OpenStudy (freckles):

a^0=1 whenever a does not equal 0

OpenStudy (anonymous):

oh oops e^0=1 so 1= 0+b b=1?

OpenStudy (freckles):

yep so we have the approximation for f(x) is y=5x+1 So now you solve 5x+1=2 and you are done.

OpenStudy (anonymous):

5x+1=2 -1 -1 -------- 5x=1 x=1/5 so exact answer is x= 0.2 ?

OpenStudy (freckles):

what that is an approximation since we approximated f(x) to be y.

OpenStudy (freckles):

oops well not what*

OpenStudy (freckles):

that is the exact approximation they are looking for

OpenStudy (anonymous):

ohh okay.. does that mean to 0.2 is the answer to this problem then? :o

OpenStudy (freckles):

yep if i have read the question correctly and by exact answer they mean the exact answer they are looking for and not the exact solution to e^(4x)+x=2

OpenStudy (anonymous):

ohh okay i see:) awesome! thank you1! :D

OpenStudy (freckles):

Np. Are you doing online homework?

OpenStudy (anonymous):

yes:/ my teacher assigned both online and written :( but i guess i get more practice! :)

OpenStudy (freckles):

fun stuff

OpenStudy (anonymous):

haha i know right? all this homework makes me jump with joy :P

OpenStudy (anonymous):

^^i wish :P